OpenStudy (anonymous):

What is the derivative of f(x)=x^(2/3) ?

OpenStudy (bibby):

Just follow the power rule\[\frac{ d }{ dx}x^2 = nx^{n-1}\]

OpenStudy (anonymous):

i haven't learned the power rule yet, only the difference quotient

OpenStudy (bibby):

This is what the power rule looks like\[\large \frac{ d }{ dx}x^n = nx^{n-1}\]

OpenStudy (anonymous):

how do you use it?

OpenStudy (bibby):

for example x^3 = 3x^(3-1) = 3x^2

OpenStudy (raden):

@bibby, you mean d/dx (x^n), right ?

OpenStudy (bibby):

lol yeah, fixed that up in the second post

OpenStudy (anonymous):

then would it be 2/3x^(-1/3) ?

OpenStudy (bibby):

exactly. You can rewrite \[\large x^{-\frac{1}{3}}\]

OpenStudy (anonymous):

1/x^1/3?

OpenStudy (raden):

where the 2 ? @bibby ?

OpenStudy (bibby):

The whole thing is multiplied by 2/3 @RadEn

OpenStudy (anonymous):

would the derivative be 2/3x^(1/3) ?

OpenStudy (bibby):

correct

OpenStudy (raden):

yeah , it is 2/3 * x^(-1/3) or 2/(3x^(1/3))

OpenStudy (anonymous):

And what type of discontinuity would this function be?

OpenStudy (anonymous):

would the function represent an infinite discontinuity?

OpenStudy (bibby):

this is what the graph looks like. I can't tell what kind of discontinuity it'd be http://www.wolframalpha.com/input/?i=2%2F%283x%5E%281%2F3%29%29